Calculate the length of a pipe that has a fundamental frequency of 997 Hz. (Take the speed of sound in air to be 343 m/s.) (a) Assume the pipe is closed at one end. m (b) Assume the pipe is open at both ends. m
Answer #1

A) IN A CLOSED PIPE L = v/4f = 343/4*997

L = 0.086 m

2) in an open pipe L = v/2f = 343/2*997 = 0.172 m